# Anisotropy Scaling Functions in Heavy-Ion Collisions: Insights into the ‘Ultra-Central Flow Puzzle’ and Constraints on Transport Coefficients and Nuclear Deformation

### v2

@@ Abstract
In summary, anisotropy scaling functions, derived from extensive measurements of $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ in Pb+Pb collisions at 5.02 and 2.76 TeV, and Xe+Xe collisions at 5.44 TeV, provide crucial insights into the ‘ultra-central flow puzzle’ and constrain transport coefficients, the eccentricity spectrum, and nucleus deformation. Consolidating diverse measurements into a single curve, these functions account for specific parameters influencing $v_{n}(p_{T},\text{cent})$, including initial-state eccentricities ($\varepsilon_{n}$), dimensionless size ($\mathbb{R}$), radial flow magnitude, the medium’s stopping power ($\hat{q}$), and its specific shear viscosity ($\eta/s$), as well as the viscous correction to the thermal distribution function ($\delta_{f}$). They yield distinct constraints for transport coefficients and nucleus deformation.
Acknowledgement
This research is supported by the US DOE under contract DE-FG02-87ER40331.A008.

### v3

@@ Abstract
In summary, anisotropy scaling functions derived from comprehensive measurements of $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ in Pb+Pb collisions at 5.02 and 2.76 TeV, Xe+Xe collisions at 5.44 TeV and Au+Au collisions at 0.2 TeV, have been instrumental in unraveling the ‘ultra-central flow puzzle’. By consolidating diverse measurements into a coherent curve, these functions illuminate key factors influencing $v_{n}(p_{T},\text{cent})$, such as initial-state eccentricities ($\varepsilon_{n}$), dimensionless size ($\mathbb{R}$), radial flow, the medium’s stopping power ($\hat{q}$), and specific shear viscosity ($\eta/s$), along with viscous corrections ($\delta_{f}$). This comprehensive approach not only deepens our understanding of the transport properties and deformation of nuclei but also provides precise constraints for modeling the quark-gluon plasma.